# ![[Uncaptioned image]](extracted/5527354/head_foot/header_bar.png) CryinGAN: Design and evaluation of point-cloud-based generative adversarial networks using disordered materials – application to LiScCl-LiCoO battery interfaces | Adrian Xiao Bin Yong, and Elif Ertekin |

### v1

@@
![[Uncaptioned image]](extracted/5527354/head_foot/header_bar.png)
CryinGAN: Design and evaluation of point-cloud-based generative adversarial networks using disordered materials – application to LiScCl-LiCoO battery interfaces |
Adrian Xiao Bin Yong, and Elif Ertekin |
Generative models have received significant attention in recent years for materials science applications, particularly in the area of inverse design for materials discovery. While current efforts have mainly focused on bulk materials with relatively small unit cells, the possibility of generative models for more complex, disordered materials would extent the current capabilities of generative modeling to a significantly wider range of materials. Generative models would also benefit from better ways to assess their performance. They are typically evaluated on the new, unverified materials being generated, which give limited metrics to evaluate how well the model has learnt from the training data. In this work, we design and evaluate models intended for disordered interface structure generation. Using a disordered LiScCl-LiCoO battery interface as a model disordered material interface, we tested different point-cloud-based generative adversarial network architectures that further include bond distance information in the discriminator, rather than only atomic coordinates. By working with a fixed material system, we evaluated the model performance through direct comparisons between training and generated structures. PointNet-based 1D convolutions, while lacking explicit knowledge of each atom’s neighbors, have good sampling sensitivity and resulted in stable training. On the other hand, graph convolutions led to discriminators with overly high sampling sensitivity, such that the models struggled to train useful generators. The best performing architecture, Crystal Interface Generative Adversarial Network (CryinGAN), uses two separate 1D convolutional discriminators, one that accepts coordinates and another that accepts bond distances as input. We demonstrate that CryinGAN is able to successfully generate low-interface-energy structures for systems with 250 atoms, in which the generated interfaces are structurally similar to the training structures. Local coordination motif analysis of the disordered region revealed that the generated structures exhibited the low-energy features of the training structures. This study highlights the capabilities of a relatively simple generative model in generating large disordered materials, and discusses the limitations of the point cloud representation. Insights are provided to help guide the development of future generative models that are useful to not just disordered, but also ordered materials. |
@@ 1 Introduction
Generative modeling has emerged as a powerful tool for tackling problems in materials science 1, 2. Initially limited to simpler molecules 3, 4 and proteins 5, generative modeling has since advanced to include inorganic materials 6, 7, 8 as well. The main interest involving generative modeling has been its potential for the inverse design of materials 9, 10, 11, in which the ultimate goal is to generate new materials with targeted properties, instead of screening known materials for desired properties. Generative models are distinguished from discriminative models, as the latter learns the conditional probability of observing a property () given a material representation (). Instead, a generative model learns the joint probability distribution p of the data that it was trained on, and samples from the distribution of structures.
@@
Most generative modeling efforts for inorganic materials 9, 11, 12 have been focused on simpler bulk crystals, comparatively less attention has been given to disordered systems, even though disordered materials are relevant to a wide range of applications 13, 14, 15. Disordered systems usually have complex and irregular structures, necessitating large atomic representations and requiring more powerful generative models than those developed for simple crystals. They include structures that completely lack crystal lattices such as amorphous materials, as well as structures beyond bulk materials such as surfaces and interfaces (refer to Fig. 1a). In direct physical modeling, disordered materials are most typically represented by large so-called “supercells” that (spuriously) introduce periodicity at larger length scales. Databases of disordered materials are growing 16, 17, 18 and they present exciting opportunities for the inverse design of metal-organic frameworks, porous amorphous materials, amorphous battery materials, etc. Beyond materials discovery, generative modeling can be used to generate amorphous structures of arbitrarily large sizes upon training on smaller samples that capture the correlation length of the material 19. This capability allows for better studies of properties that are sensitive to size effects such as thermal conductivity and mechanical properties. Generative modeling can also be used to optimize atomic structures to match experimental observations 20, in which the structures being optimized are often disordered 21, 22. However, the capabilities of generative modeling for disordered systems remain largely limited (e.g., generation of 2D morphology 19 instead of actual atomic structures), and generative models have been reported to fail for large systems 11, 23. To reap the benefits of generative modeling for disordered systems, better generative models need to be developed and tested on disordered systems.
@@
When building a generative model, two major design decisions are the type of generative model and the material representation (i.e., the input used to describe the material). The compatibility of these two choices is important as well. The types of generative models that have been used for materials include variational autoencoders (VAEs) 7, 11, 9, 6, generative adversarial networks (GANs) 24, 8, 25, 10, diffusion models 26, 27, 28, 29, and language models 30, 31, 32. Generative models were initially developed with two main types of material representations: (1) voxels 6, 7, 11, 12 and (2) point clouds 8, 25, 9, 10. Voxel representation treats cells as voxelized density maps. New structures are generated as density maps that are subsequently reconstructed back into atomic coordinates. One disadvantage of this representation is that it is memory intensive, resulting in limitations in the voxelization resolution and thus number of atoms (e.g., Court et al. 11 restricted the number of atoms to 40 atoms per cell). Reconstruction issues were also reported 11 for non-cubic cells. These considerations make voxel representations less suitable for disordered materials. On the other hand, point clouds directly represent structures using their atomic coordinates and lattice parameters, circumventing the memory and reconstruction issues of the voxel representation. The simple representation makes them a highly scalable and better choice for disordered materials. However, the design of point cloud architectures that are symmetry-invariant is not trivial, as the commonly used PointNet architecture 33 does not include the desired symmetry invariances. More recently, other representations such as graphs 26, 29, coordinate-based representations (e.g., UniMat 28, CrysTens 27), and text-based representations 30, 31, 32 have also been explored.
@@
To compare different generative models and make design choices, it is necessary to sufficiently evaluate the generated structures for their validity. Training generative models for materials discovery inherently makes the evaluation of the models’ performance difficult. In more conventional problems such as image generation or speech synthesis, it is relatively easy to discern if the model has learned from the training data to generate realistic images or speech. However, it is much more difficult to determine if a newly generated material is realistic, or if the model is simply generating noise. With only limited and subjective metrics describing material validity, it is difficult to evaluate which model design choices yield better performance. One approach to circumventing this issue is to train the model on a fixed set of materials (e.g., perovskites). Restricting the material space allows for easier determination of whether the correct structure is being generated, and direct comparison of the properties between the generated and training structures can be performed. In this scenario, however, the materials on which the model is trained should be sufficiently challenging to provide meaningful evaluation of the model’s performance. In this regard, disordered materials are good candidates for the task, given that generative models can fail to generate even a single type of disordered material (e.g., amorphous silicon 23).
In this work, we demonstrate a design and evaluation process for generative models for a fixed set of disordered materials. We show how this process is valuable to improving generative models in general (and not just for disordered materials). We chose the commonly used point cloud representation for this work, and tested multiple different GAN architectures where we included bond distance information explicitly in the GANs. The performances of these GANs were evaluated using a LiScCl(100)-LiCoO(110) disordered battery interface system (see Fig. 1b). The GANs were tasked with generating interface structures with low interface energy. Comparison of GAN performances reveal the architecture that is most effective at accomplishing this goal, as well as reasons why the other architectures do not perform as well. We demonstrate that the best architecture, Crystal Interface Generative Adversarial Network (CryinGAN), can generate large disordered interfaces with low interface energy, and similar structural features to the training structures. Through this work, we provide insights regarding generative model design and promising directions to pursue for future models of disordered materials.
@@ 2 Results and discussion
2.1 Interface disorder mechanism
The disordered interface considered in this work is a solid-state battery interface between the LiCoO (LCO) cathode and the LiScCl (LSC) solid electrolyte. LCO is one of the most commonly used cathode materials in commercial Li-ion batteries 34. LSC 35 belongs to the class of halide solid electrolytes, which can achieve both high ionic conductivity and high-voltage stability, enabling the use of high-voltage cathodes in all-solid-state batteries 36, 37. In previous work 38 focused on the interface between LCO and a different halide solid electrolyte, LiYCl (LYC), we found the formation of a highly disordered LCO-LYC interfacial structure. This disordered interfacial structure was similarly observed for the calculated LCO-LSC interfaces considered in this work, despite the different compositions and crystal structures of the halide solid electrolytes.
Disordered interfaces have been observed experimentally across diverse material systems. They arise for a variety of reasons including elemental segregation to the interface 39, 40, intermixing across the interface 41, and ion irradiation 42. For the material system of this study, we found that chlorides have an innate tendency to form disordered interfaces with oxides. In this subsection, we provide an explanation of the interface disorder mechanism for completeness, before presenting on generative modeling of the LCO-LSC interfaces. While many heterointerfaces can adopt regular epitaxial registries connecting two materials with well-defined crystalline orientations, other heterointerfaces show irregular, disordered interfacial patterns. As shown in Fig. 2, using binary instead of ternary materials, we found that disordered interfacial structures are obtained with only the combination of LiO and LiCl. We suggest two interdependent reasons for their occurrence. The first reason is the bond formation between O (in the oxide) and Li (in the chloride). Fig. 2a shows the interface structure of LiO(100)-LiCl(100) for LiO(100) slabs that are either Li-terminated or O-terminated. When LiO(100) is Li-terminated, there is minimal rearrangement at the interface with only some Li-Cl bond formation. In contrast, when LiO(100) is O-terminated, significant atomic rearrangement is observed with a combination of Li-O and Li-Cl bond formation. The stronger interaction between O and Li is likely due to the higher charge density of O compared to Cl. However, Li-O bond formation alone does not lead to highly disordered interfaces, since such degree of disorder is not observed in other O-containing interfaces such as oxide-oxide and oxide-sulfide interfaces 43, 44.
The second reason that the chlorides can form disordered interfaces is due to their intermediate mechanical stiffness. To form bonds with the O in the oxide slab, the other slab must distort its crystal structure to align its atoms with O, and the stiffness of the material determines the extent of the distortion allowed. Fig. 2b shows the interface structure between LiO and materials in increasing order of mechanical stiffness (Li, LiCl, and MgS). MgS has the same rock salt crystal structure as LiCl but with higher stiffness 45, 46. Due to its higher stiffness, it is unable to distort significantly to bond with O. On the other hand, LiCl is soft enough that it can distort itself, whilst maintaining some of its structure, resulting in a more distorted interface structure. For a very soft material like Li metal 47, the atoms can rearrange themselves relatively freely to bond with O, so large voids such as those observed for LiCl are less likely to form. The intermediate stiffness of the chlorides, combined with bond formation with O, results in the chloride solid electrolytes’ ability to form highly disordered interface structures with the oxide cathode. The disordered nature of the LCO-LSC interfaces, as shown in Fig. 1b, is therefore understood from the two considerations above.
2.2 Dataset generation
@@
To train the interface GAN models, we first generated a dataset of LSC-LCO interface structures. A brief overview of the methodology is outlined here, further details can be found in the Methods section. LiScCl(100)-LiCoO(110) interfaces (see Fig. 1b for reference) were generated by randomly generating 3 formula units of LSC atoms in the interface region between the LSC and LCO slabs. For each structure, the thickness of the interface region was randomly selected between 4 and 6 Å, and a random lateral displacement was applied to the LSC slab (translation allowed along the full range of both lateral directions). The randomly generated structures were relaxed using density functional theory (DFT) calculations.

### v2

# ![[Uncaptioned image]](extracted/5729874/head_foot/header_bar.png) Dismai-Bench: Benchmarking and designing generative models using disordered materials and interfaces | | Adrian Xiao Bin Yong,∗ab Tianyu Su,ab and Elif Ertekin∗bc | |

@@
![[Uncaptioned image]](extracted/5729874/head_foot/header_bar.png)
Dismai-Bench: Benchmarking and designing generative models using disordered materials and interfaces |
| Adrian Xiao Bin Yong,∗ab Tianyu Su,ab and Elif Ertekin∗bc | |
Generative models have received significant attention in recent years for materials science applications, particularly in the area of inverse design for materials discovery. However, these models are usually assessed based on newly generated, unverified materials, using heuristic metrics such as charge neutrality, which provide a narrow evaluation of a model’s performance. Also, current efforts for inorganic materials have predominantly focused on small, periodic crystals ( 20 atoms), even though the capability to generate large, more intricate and disordered structures would expand the applicability of generative modeling to a broader spectrum of materials. In this work, we present the Disordered Materials & Interfaces Benchmark (Dismai-Bench), a generative model benchmark that uses datasets of disordered alloys, interfaces, and amorphous silicon (256-264 atoms per structure). Models are trained on each dataset independently, and evaluated through direct structural comparisons between training and generated structures. Such comparisons are only possible because the material system of each training dataset is fixed. Benchmarking was performed on two graph diffusion models and two (coordinate-based) U-Net diffusion models. The graph models were found to significantly outperform the U-Net models due to the higher expressive power of graphs. While noise in the less expressive models can assist in discovering materials by facilitating exploration beyond the training distribution, these models face significant challenges when confronted with more complex structures. To further demonstrate the benefits of this benchmarking in the development process of a generative model, we considered the case of developing a point-cloud-based generative adversarial network (GAN) to generate low-energy disordered interfaces. We tested different GAN architectures and identified reasons for good/poor performance. We show that the best performing architecture, CryinGAN, outperforms the U-Net models, and is competitive against the graph models despite its lack of invariances and weaker expressive power. This work provides a new framework and insights to guide the development of future generative models, whether for ordered or disordered materials. |
@@ 1 Introduction
Generative modeling has emerged as a powerful tool for tackling problems in materials science 1, 2. Initially limited to simpler molecules 3, 4 and proteins 5, generative modeling has since advanced to include inorganic materials 6, 7, 8 as well. The primary interest in generative modeling has been its promise for inverse materials design 9, 10, 11, where the primary objective is to create new materials tailored to specific properties rather than screening known materials for desired characteristics. Generative models are distinguished from discriminative models, as the latter learns the conditional probability of observing a property () given a material representation (). Instead, a generative model learns the joint probability distribution of the data that it was trained on, and samples from the distribution of structures.
@@
While generative modeling efforts for inorganic materials 9, 11, 12 have primarily centered around simpler bulk crystals, there has been comparatively less emphasis on disordered systems, despite their relevance across a wide spectrum of applications 13, 14, 15. Disordered systems usually have complex and irregular structures, necessitating large atomic representations and requiring more powerful generative models than those developed for simple crystals. They include structures that completely lack crystal lattices such as amorphous materials, as well as structures beyond bulk materials such as surfaces and interfaces. In direct physical modeling, disordered materials are most typically represented by large so-called “supercells” that (spuriously) introduce periodicity at larger length scales. Databases of disordered materials are growing 16, 17, 18, offering compelling prospects for the inverse design of metal-organic frameworks, porous amorphous materials, amorphous battery materials, and more. Beyond materials discovery, generative modeling can be used to generate amorphous structures of arbitrarily large sizes upon training on smaller samples that capture material correlation lengths 19. This capability enables more thorough investigations into properties that are influenced by size effects, such as thermal conductivity and mechanical properties. Generative modeling can also be used to refine atomic structures to align with experimental observations 20, typically focusing on the refinement of disordered structures 21, 22. Yet, the application of generative modeling to disordered systems remains limited, such as generating 2D morphology rather than precise atomic structures 19. Moreover, generative models have been reported to fail when applied to large systems 11, 23. To reap the benefits of generative modeling for disordered systems, better generative models need to be developed and evaluated on disordered systems.
@@
When building a generative model, two major design decisions are the type of generative model and the material representation (i.e., the input used to describe the material). The compatibility of these two choices is important as well. The types of generative models that have been used for materials include variational autoencoders (VAEs) 7, 11, 9, 6, generative adversarial networks (GANs) 24, 8, 25, 10, diffusion models 26, 27, 28, 29, 30, and language models 31, 32, 33. Generative models were initially developed with two main types of material representations: (1) voxels 6, 7, 11, 12 and (2) point clouds 8, 25, 9, 10. Voxel representation is memory intensive, resulting in limitations in the voxelization resolution and thus number of atoms (e.g., Court et al. 11 restricted the number of atoms to 40 atoms per cell). Reconstruction issues were also reported 11 for non-cubic cells. On the other hand, point clouds directly represent structures using their atomic coordinates and lattice parameters, making them highly scalable with the number of atoms. However, the design of point cloud architectures that are symmetry-invariant is not trivial, as the commonly used PointNet architecture 34 does not include the desired symmetry invariances. More recently, other representations such as graphs 26, 30, coordinate-based representations (e.g., UniMat 29, CrysTens 28), and text-based representations 31, 32, 33 have also been explored. Graphs are particularly attractive due to their symmetry invariances and strong expressive power, capturing both geometrical features and neighbor information. However, graph convolutions become computationally and memory intensive as the number of atoms increases.
@@
To compare different generative models and make design choices, it is necessary to sufficiently evaluate the generated structures for their validity. Training generative models for materials discovery inherently makes the evaluation of the models’ performance difficult. In more conventional problems such as image generation or speech synthesis, it is relatively easy to discern if the model has learned to generate realistic images or speech from the training data. However, it is much more difficult to determine if a newly generated material is realistic, or if the model is simply generating noise. Recent generative models 26, 27, 29, 33 have relied on limited and heuristic metrics (e.g., charge neutrality, material space coverage) to evaluate and compare between models, making it difficult to meaningfully assess model performance. One approach to circumventing the issues of evaluating on new, unknown materials is to instead train the model on a fixed set of materials (e.g., perovskites). Restricting the material space allows for easier determination of whether the correct structure is being generated, and direct comparison of the properties between the generated and training structures can be performed. In this scenario, however, the materials on which the model is trained should be sufficiently challenging to provide meaningful evaluation of the model’s performance. In this regard, disordered materials are good candidates for the task, given that generative models can fail to generate even a single type of disordered material (e.g., amorphous silicon 23).
In this work, we present the Disordered Materials & Interfaces Benchmark (Dismai-Bench), a generative model benchmark that uses datasets of an Fe60Ni20Cr20 austenitic stainless steel, a disordered Li3ScCl6(100)-LiCoO2(110) battery interface, and amorphous silicon. Dismai-Bench evaluates generative models on a wide range of material disorder ranging from structural to configurational (see Fig. 1). Structural disorder increases from left to right, and configurational disorder increases from right to left, in Fig. 1. The composition of each dataset is fixed, and each structure has 256-264 atoms. We selected four recent diffusion models to be benchmarked on Dismai-Bench, including two models that use graph representations (CDVAE 26 & DiffCSP 27) and two models that use coordinate-based representations (CrysTens 28 & UniMat 29). The models were trained on one dataset at a time, and the generated structures were compared with the training structures to obtain structural similarity metrics. These metrics quantify the model’s ability to learn complex structural patterns found in disordered materials. We show that the graph models outperform the coordinate-based models due to the higher expressive power of graphs. The success of the less expressive models in materials discovery 29, 33 suggest that noisy models are better for discovering small crystals, but face challenges when tasked with generating larger, more complex structures.
To demonstrate the application of Dismai-Bench in the development of a generative model, we further considered the design of a GAN to generate low-interface-energy Li3ScCl6(100)-LiCoO2(110) interface structures. We chose the simple point cloud representation, and tested multiple different GAN architectures for which we included bond distance information explicitly in the GANs, instead of just atomic coordinates. Direct comparison between the generated and training structures identified the architecture that best achieved the intended goal, along with explanations for why the other architectures were less successful. We demonstrate that the best architecture, Crystal Interface Generative Adversarial Network (CryinGAN), can generate the disordered interfaces with low interface energy, and similar structural features to the training structures. Despite its design simplicity, CryinGAN outperforms the more recent coordinate-based diffusion models on Dismai-Bench. It does not outperform the graph diffusion models across all datasets, however, possibly as a result of its weaker expressive power and lack of invariances. Through this work, we present a novel framework for conducting meaningful comparisons between models, providing valuable insights into model weaknesses and failures to inform the design of future generative models.
@@ 2 Results and discussion
2.1 Datasets and interatomic potentials
An overview of the dataset curation is outlined here; further details are found in the Methods section. A total of six datasets are used in Dismai-Bench, consisting of four alloy datasets, one interface dataset, and one amorphous silicon dataset. Each dataset contains a total of 1,500 structures, split into 80 % training and 20 % validation data. Test sets are not needed since model performance is measured using the benchmark metrics.
2.1.1 Fe60Ni20Cr20 austenitic stainless steel
The stainless steel datasets consist of face-centered cubic (FCC) crystals that are structurally simple, but configurationally complex (refer to Fig. 1). Atoms of various species occupy the lattice sites with different ordering tendencies. The generative models are challenged with generating structures that not only have well-defined FCC lattices, but also the correct degrees of ordering. The stainless steel datasets were created using a cluster expansion Monte Carlo (CEMC) approach 35, 36. The cluster expansion (CE) potential was adapted from ref. 36 , where pair interactions up to the 7 neighbor shell were included. The composition of the structures is Fe60Ni20Cr20, and each structure contains 256 atoms. Monte Carlo (MC) simulations were carried out in the canonical ensemble at 300 K and 1,500 K, such as to obtain datasets with different degrees of short-range order (SRO). The SRO is quantified using the Warren-Cowley SRO parameter 37,
where is the probability of finding pairs in the -th neighbor shell, and = / is the conditional probability of finding atom in the -th coordination shell of atom . and are the concentration of and atoms respectively. SRO parameter indicates zero correlation between atoms (as in a random solution), while indicates an attractive interaction and indicates a repulsive interaction. Note that the and atoms can be of the same atomic species.
The SRO distributions of the 300 K and 1500 K alloy training datasets are shown in ESI Fig. S1 for the 1st and 2nd nearest neighbor interactions. The 300 K dataset shows more prominent SRO than the 1500 K dataset. The SRO parameter tends to distribute away from zero at 300 K, and consistently distribute near zero at 1500 K. We also created two additional datasets by filtering the CEMC-generated structures such that the SRO parameters have narrow distributions within 0.1 of the average values (see ESI Fig. S2). Sufficiently large number of structures were generated to obtain 1,500 structures for each dataset. We refer to the unfiltered and filtered datasets as the wide SRO and narrow SRO datasets respectively. This SRO filtering was performed to enable comparison of generative model performance when trained on structures with noisy SRO distribution (wide SRO dataset) and structures with less noisy SRO distribution (narrow SRO dataset).
2.1.2 Amorphous silicon
Amorphous silicon can be thought of as the polar opposite of the FCC alloys. Amorphous silicon consists of a single atomic species only, but completely lack ordering in the form of a crystalline lattice. The generative models are not assessed on any ability to learn ordering relationships between different atomic species. Instead, they are assessed on their abilities to learn the complex structural patterns found in amorphous silicon, such as near-tetrahedral local environments and pair distribution functions. The amorphous silicon dataset was adapted from ref. 38. The original data consists of a 100,000-atom amorphous silicon structure generated through melt-quench molecular dynamics simulation 38. The structure was sliced into smaller blocks with lattice parameters corresponding to 256-atom amorphous silicon structures. The blocks were sliced at different locations to obtain a total of 1,500 blocks. Blocks with < 256 atoms had atoms added at random to low density regions, and blocks with > 256 atoms had atoms removed at random from high density regions, so that all blocks have 256 atoms. The 1,500 structures were relaxed using a pre-trained SOAP-GAP 39 machine learning interatomic potential for Si. The resulting structures have a higher concentration of defects compared to the original 100,000-atom structure (refer to ESI Fig. S3) due to the slicing and atom addition/removal, but the Si coordination geometry remains predominantly tetrahedral.
2.1.3 Li3ScCl6(100)-LiCoO2(110) battery interface
The disordered interface dataset assesses the generative models on structures that exhibit a mixture of structural and configurational disorder (refer to Fig. 1). Atoms in the disordered interface region are not arranged in well-defined lattices, and coordinate with each other in a range of motifs. The models have to learn to generate these complex heterogeneous interfaces, where they need to construct the crystalline slabs and disordered interface region correctly. The interface considered in this work is a solid-state battery interface between the LiCoO2 (LCO) cathode and the Li3ScCl6 (LSC) solid electrolyte. LCO is one of the most commonly used cathode materials in commercial Li-ion batteries 40. LSC 41 belongs to the class of halide solid electrolytes, which can achieve both high ionic conductivity and high-voltage stability, enabling the use of high-voltage cathodes in all-solid-state batteries 42, 43. In previous work 44, a similar disordered interface structure was observed between LCO and a different halide solid electrolyte, Li3YCl6, despite the different compositions and crystal structures of the halide solid electrolytes. Disordered interfaces have been observed experimentally across diverse material systems. They arise for a variety of reasons including elemental segregation to the interface 45, 46, intermixing across the interface 47, and ion irradiation 48. For the oxide-chloride interface system, we found that chlorides have an innate tendency to form disordered interfaces with oxides; further details are documented in ESI Supplementary Note 1.
@@ 2.1.3 Li3ScCl6(100)-LiCoO2(110) battery interface
The Li3ScCl6(100)-LiCoO2(110) interface dataset was created by generating random interface structures and relaxing them. Each structure was first constructed by randomly generating 3 formula units of LSC atoms in the interface region between the LSC and LCO slabs. For each structure, the thickness of the interface region was randomly selected between 4 and 6 Å, and a random lateral displacement was applied to the LSC slab (translation allowed along the full range of both lateral directions). The randomly generated structures were relaxed using density functional theory (DFT) calculations.